### v1

@@ 3.2 Features of this platform
  • Unification: The platform can collect, store, and process all types of data from various sources, regardless of volume, velocity, and variety, thanks to its modern open data architecture. This feature contributes to:
    • The development of batch and real-time processing for network applications.
    • Enabling data analysts to process and access data for advanced analytics tasks.
    • Supporting the development and deployment of AI models, paving the way for advanced federated learning in multi-vendor networks.
    • Handling operational needs ranging from reporting of complex processed metrics, performance, or fault events to interactive visualization with these dashboards and data cubes.
  • Cost-effective: Built on commodity servers, the platform leverages Kubernetes (K8s) container orchestration and other tools to simplify management, reducing the resources required. On the other hand, Using open data architecture and data lakehouse technology reduces costs by consolidating data, optimizing query performance, reducing bandwidth usage, leveraging scalable cloud storage, utilizing cost-effective open-source tools, and improving collaboration and usability.
@@
  • Scalability: As the number of O-RAN networks deployed increases the platform’s scalability is crucial. It has been thoroughly evaluated to ensure it meets the increased demands for AI, data processing, and storage.
3.3 Platform core components:
Figure 5: Boldyn Networks cloud-native data analytics core components. This figure shows the main components related to infrastructure, Operating systems, k8s management and data core components.
The first tier, following a bottom-up order, contains the K8s infrastructure and its automation toolset. we utilized Terraform for defining and provisioning our infrastructure as code enabling efficient management and automation of infrastructure resources that run the operating system (second tier of the architecture) Talos OS. Talos OS is a modern, Linux-based operating system that is specifically designed for K8s. It provides a secure, minimal, and immutable platform to enhance security. Its major feature is that it automates and simplifies the management and operations of K8s clusters. On top of this, K8s is installed to deploy the microservices and applications that implement K8s cluster management functions and data management applicational system (this being the third tier). The following applications and services are implemented in the third tier:
3.3.1 Data management applicational systems
  • Apache NiFi is the core of the Acquisition and Mediation layer and is used to automate and manage the flow of data between systems where the data sources are located (across the multiple platforms and their components) and other applications and systems of the data management platform that consume these data. Therefore, it is the main application used to implement the collection and mediation layer of the platform.
  • Apache Kafka is the core streaming system and is used to transfer data in real-time from the source to the processing layer.
  • In the storage layer, we use the object storage Minio which provides an S3-like API. Depending on the data use case, we store the data using one of the data storing techniques like Hudi, Deltalake, or Iceberg open tables, or even in its raw status.
  • The data is later accessible by Apache Spark to perform complex big data processing on data from Streaming, data lakehouse, databases, or from all. Trino is, on the other hand, a virtualization system and it’s used to perform SQL-like queries on any dataset available anywhere in the platform.
  • In the application layer, we use Python, Jupyter Notebook, and MLflow to train, test, and validate the AI module before publishing its micro-service deployment in the processing layer. We also use DBT for version-controlled analytics workflows.
  • In the control and management phase, we use Apicorio as a schema registry, Hive Metastore to store data catalogues and metadata, and we also use PostgresDB to store the network topology, rules, and alarm triggers.
  • The last phase in the data is Elasticsearch and Kibana where we visualize different data from different sources in addition to the applied AI actions and results in one dashboard that serves as the single pane of glass in the platform.
3.3.2 K8s cluster management tools
There are other systems mentioned in the management toolbar. These tools are used to support the data tools in the data layers.
  • Longhorn & Rancher: for k8s and cloud-native storage (data layer for applications) management.
  • Prometheus and Grafana: for platform monitoring and alerting.
  • NeuVector and Kube-bench: for security and vulnerability checks.
  • Velero and Fluentd: for logs and backup management.
  • Argo CD and GitHub Actions: for CI/CD and GitOps.
3.3.3 CI/CD pipeline
The development of data pipelines on this architecture can be a complex, tedious, and error-prone process if done manually. A set of tools and systems are used to automate the development and deployment of these pipelines. GitOps methods, such as CI/CD, are used to simplify, automate and manage this process.
Fig. 6 shows an example of one CI/CD process implemented to develop and deploy the data pipeline of one use case addressed by the platform, providing details about every data engineering and AI model preparation task involved.
The pipeline CI/CD cycle starts by developing the source code and publishing it onto the GitHub repository, which triggers a set of automated workflows that checks the quality of the code and performs security scans. In the next step, the code is built into Docker images following a process known as containerization. All the Docker images are then submitted to a process of security vulnerability scanning, before being stored in the Docker registry. The last stage of the CI process is completed when all the deployment manifests are updated with the new Docker image. The final stage of the cycle is the CD process, which uses ArgoCD to automate the process of fetching the latest changes in the deployment manifests and deploying the new application.
Figure 6: Automation pipelines for the AI and related data processes in the platform.
4 The problem statement and the proposed solution
4.1 Problem statement
Fig. 7 illustrates the use case of deploying a commercial O-RAN-based MPN in an offshore location, serving the UEs that are located in vessels that navigate around the windfarm attending to wind turbines for infrastructure operational and maintenance activities. These vessels spend most of the time navigating close to the windfarm and carry tens of people who work at sea for periods of 15 days. Therefore, these people rely on this connectivity to do their work, to communicate with their colleagues working onshore, their families, and for their entertainment. The connectivity provided by this network also supports business and operational critical processes to the organizations responsible for operating and maintaining the network, across an area of around 300 Km2. Many factors such as weather conditions, sea conditions, distance to the site, and MPN operational faults may affect the quality of the connectivity service as experienced by the end-users, which might impact the: ability of running business and operational critical processes work according to the requirements or the ability of people working productively in their floating offices. Despite these challenges, the MPN’s network operator is responsible for managing, operating, and maintaining the network’s performance as per the contracted service level agreement.
Figure 7: Illustration for the real-time offshore 5G network deployment using ORAN and challenges in managing coverage in the sea.
The problem addressed by the solution described in this paper occurred in an MPN service deployed by Boldyn Networks in a windfarm located in the North Sea. This network is composed of 3-cellular sites of 3 sectors each and one cellular carrier per sector, totaling 9 macro cells that cover the whole extension of the windfarm. Each cell operates in the LTE B3 and implements a 20MHz channel. The cellular sites are installed across the windfarm in three different turbines, where commercial-of-the-shelf (COTS) servers are used to run the containerized O-DU, and three O-RUs are installed to implement a sector each. The O-DUs deployed offshore connect, via dark-fibre light-up using long-haul SFPs, to an infrastructure of COTS servers that run the O-CU container. On the other hand, each relevant vessel’s network is composed of a Wi-Fi network that provides IP connectivity across the vessel. This Wi-Fi network connects to Boldyn’s customer premise equipment (CPE) that acts like an LTE broadband router back-hauling the vessels’ IP traffic to connect to the internet and the MPN customer’s enterprise network. This CPE is installed inside the vessel, near the bridge, and it’s connected to external 4x4 multiple-input-multiple-output (MIMO) antennae that increase the coverage and capacity of the network. Each CPE is dual-modem capable for connection resilience and traffic load balancing reasons and one MPN SIM is inserted in each modem. The problem resides in the connection recovery procedure implemented by the CPE when one of the modems drops its link to the LTE network. This procedure takes 5 minutes to re-establish an LTE connection to the macro network which is a long period of time and in some conditions increases the likelihood of disconnections occurring in both modems at the same time. In studying the problem, it was found the following:
  • It could be observed drift in the behavior of radio link performance, between the two modems of the CPE, verified by monitoring indicators such as latency and throughput even if most of the time these were connected to the same cell and benefiting from similar radio conditions.
  • Updating the configuration of the modem on the CPE via its API, forced the modem connection to restart and established a new radio link improving the performance of the modem in all the indicators.
  • The process of restarting the modem is much quicker to implement than the process of re-connection in case of radio link drop and has the additional advantage that can be done pre-emptively whilst the other modem is showing good performance.
  • The automation of this process of analysis and decision-making could reduce significantly the number of disconnections thus having a great impact on the quality of experience (QoE) of the users being served by the network system.
The automation of this process enabled us to develop a self-healing mechanism that implements an ML model, for the detection of the event of this type of performance anomaly in the modem, trained on available historical performance data merged with data that showed the timestamps where actions were taken to resolve the issue, which resulted in a set of labeled data emphasizing the relationship between patterns of behavior of the relevant performance indicators and the need for decision-making action. Some of the relevant performance indicators were reference signal received power (RSRP), reference signal received quality (RSRQ), IP packet data latency, timestamp, location, and cell to which the modem is connected to, and others from different parts of the network.
The process of collecting the data from the CPE device via the API every 5 seconds, preparing and enriching the data to stream this data to the application layer to be consumed by the ML model for online decision-making and direct provisioning, benefited from all the features and capabilities offered by the data management platform.
4.2 The proposed solution
@@
Network anomaly detection or prediction is a complicated task. Anomalies are referred to as patterns in data that do not conform to a well-defined characteristic of normal patterns. Anomalies can be classified into three types: (1) point anomaly can be considered as a particular data instance deviation from the normal pattern of the dataset; (2) contextual anomaly is defined when a data instance behaves anomalously in a particular context; (3) collective anomaly happens when a collection of similar data instances behave anomalously with respect to the entire dataset, the group of data instances is termed a collective anomaly [27]. Existing anomaly detection methods include NN, support vector machine or rule-based classification, statistical signal process method and clustering. In this case, the prediction task is based on the historical records of network states and automated actions, so it can be regarded as a point anomaly issue, and NN is the preferred method. Moreover, to exploit the temporal correlations of the historical samples, the LSTM is the selected learning model.
4.3 Data source
This case study has benefited from utilizing data from a real-life network based on O-RAN standards. The O-CU, O-DU, and O-RU provide FCAPS data, which describes network performance and operational behavior. This FCAPS data is a fundamental building block for the AI model’s development cycle, playing a crucial role in the training and validation process. Once deployed, the AI model uses this data for inference. Another important data source comes from the CPE equipment in the vessel, which provides measurements of the modem’s performance and its location as a time series, mapped with the FCAPS information collected from the O-RAN network.
The CPEs are at sea, and their location is monitored continuously. A couple of months of this data is shown in Fig. 8111The latitude and longitude values are removed for the consideration of the users privacy.. This data shows records of vessels connected to the network even when they are outside the coverage area. It contains many outliers that need to be removed, specifically those indicating locations far from the nearest site.
@@ 4.3 Data source
Figure 8: Visualisation for the location of the vessel from the data used in the AI model layered on top of the predicted network coverage.
4.4 LSTM model design
Table 1: Template for some of the exported network state data
WAN ID Carrier LTE-RSRP LTE-SINR LTE-RSRQ Latency Time Latitude Longitude
@@ 4.4 LSTM model design
The development of the LSTM model adopts the manner of offline development and online deployment. The historical data stored by the platform is used for the model’s training. After the trained model meets the training criteria, the natural model will be deployed in the further step. Historical data includes two parts, the network state data and corresponding actions. Actions here are operations recorded by the engineers manually when the QoS is lower than the threshold. Some of the network state data is exported using the template shown in table 1, where WAN ID indicates the ID of the devices that need an action; the carrier is the carrier frequency; LTE-RSRP, LTE-SINR, LTE-RSRQ, Latency, Latitude and Longitude are UE related information; Time refers the timestamp for that record.
@@
The next anomaly must be forecasted according to the hidden pattern of historical network state records, so a reasonable solution is to formalize a sequence prediction task, where multiple discrete records on sequential timestamps will be cascaded as one training sequence. Then these training sequences are labelled by using manual operations. The training sequence is labeled as ‘0’ if there is an action needed, otherwise, the label is ‘1’.
For the features in table 1, the WAN ID, LTE-RSRP, LTE-SINR, LTE-RSRQ, and Latency indicators are some of the taken features for training, while removing the carrier, latitude, and longitude because they are unlikely correlated with the occurrence of the anomaly. The recording interval of raw data is 5 seconds. Two months of records were stored and then processed in the cloud-native data platform. The data pre-processing process is illustrated in Fig. 9.
@@
First, for the given raw dataset, there are some records missing. So, the process needs to pad these missing values and remove the records that are obviously out of the normal range. Then as a step of de-noising, the data is downsampled to $1/N$ of the initial one. Then assemble every $L$ sample as one training sequence. Lastly, the training sequence is labeled. It is worth mentioning that the criteria of labeling, are the manual actions taken by the engineer, which may lag over the first occurrence of the anomaly of $5$ to $20$ minutes. So, $10$ minutes is taken as the average delay applied to data labeling. The training sequence will be a matrix with size $1\times 5\times L$, and the labels will be binary values. In this use case, after validation, $N=2$ and $L=6$ are reasonable options.
@@
The architecture of the NN model is depicted in Fig. 10, wherein the LSTM module inherits from the LSTM function in Pytorch and the input size equals $1\times 5\times 6$, hidden size equals 2. Two linear layers follow the LSTM module, and the output is the indicator of taking action or not. The last layer adopts the activation function ’sigmoid’.
@@
For the model’s training, the training/validation dataset is divided according to the 80%/20% rule. The loss function is the binary cross entropy (BEC) [28]. The training is on the two RTX 2080Ti GPUs, the batch size is $2048$ and Adam is taken as the optimizer. In the batch sampling process, the weighted batch sampler is used because this dataset is an imbalanced dataset, which means that labels ’$1$’s are far more than labels ’00’s. The above training dataset is used to train the LSTM model, and the test set is used to evaluate the trained model. In the training stage, the training data is shuffled before batch sampling. Fig. 11 shows the model’s accuracy plots in the training and test. The convergence is reached after $100$ epochs.
@@
4.5 Model deployment and validation
The real-world deployment is relatively straightforward using the Boldyn data analytic platform. As illustrated in Fig. 4, the development happens in its application layer. In validation and deployment, the model is dockerized to an image and serves as a part of the DevOps pipeline and is hosted directly in the processing layer. This microservice continuously receives the real-time data and feeds it to the LSTM model. The generated actions will trigger the change in the networks. After deploying the model in production the complaints and network dis-connectivity decreased significantly to a very acceptable limit. Table 2 shows the improvement in the service performance and that validates the model in production.

### v2

@@ 4.5 Features of this platform
  • Unification: The platform can collect, store, and process all types of data from various sources, regardless of volume, velocity, and variety, due to its modern open data architecture. It enables batch and real-time processing, allows analysts to perform advanced data exploration, supports the development and deployment of AI models including federated learning, and provides comprehensive tools for reporting and interactive visualization across complex network environments.
  • Cost-effective: Built on commodity servers, the platform leverages Kubernetes (K8s) container orchestration and other tools to simplify management and reduce the required resources. In addition, using open data architecture and data lakehouse technologies helps consolidate data, optimize query performance, reduce bandwidth usage, and leverage open-source tools, improving collaboration and overall usability.
@@
  • Scalability: As more O-RAN networks are deployed, the ability to scale becomes critical. The platform has been thoroughly evaluated to ensure it can handle growing demands in AI workloads, data processing, and storage capacity.
5 The proposed solution
For the specific business problem introduced in Sec. 3, this section will discuss the anomaly types and model selection, data sources, LSTM model design, model deployment procedure, and validation results.
5.1 Anomaly types and model selection
@@
Network anomaly detection or prediction is a complicated task. Anomalies are referred to as patterns in data that do not conform to a well-defined characteristic of normal patterns. Anomalies can be classified into three types: (1) point anomaly can be considered as a particular data instance deviation from the normal pattern of the dataset; (2) contextual anomaly is defined when a data instance behaves anomalously in a particular context; (3) collective anomaly happens when a collection of similar data instances behave anomalously with respect to the entire dataset, the group of data instances is termed a collective anomaly [33]. Existing anomaly detection methods include NN, support vector machine or rule-based classification, statistical signal process method and clustering. In the case of the offshore O-RAN deployment studied in this work, the prediction task involves detecting abnormal behavior in modem connectivity based on historical telemetry and operational data. These anomalies are best characterized as point anomalies, as they are typically reflected in sudden shifts in performance indicators such as RSRP, RSRQ, latency, and signal quality. Given the temporal nature of these patterns and the necessity of capturing sequential dependencies in the data, LSTM networks were selected as the most suitable model. LSTM architectures are well-suited for modeling time-series data with complex temporal dynamics, allowing the system to identify early signs of degradation that precede connectivity loss. The choice of LSTM aligns with the objective of the platform to provide real-time, proactive anomaly detection within a streaming data environment. This integration enables the model to not only identify anomalous conditions but also trigger automated corrective actions with minimal latency, improving the resilience of the O-RAN deployment.
5.2 Data specification
The dataset used in this study was collected from a production deployment of a private offshore mobile network based on a O-RAN architecture. This network serves vessels operating within a windfarm environment and was instrumented with custom telemetry agents and monitoring tools to support continuous observation. Data was captured across both network-side components, like the O-CU, O-DU, and O-RU, and from onboard CPE installed on the vessels themselves. Together, these sources provided a comprehensive view of both infrastructure performance and user-side experience in a complex and mobile offshore setting. Network-side data was in line with 3GPP specifications. This included a range of FCAPS metrics such as latency, throughput, availability, signal quality indicators, Handover, fault and alarm records, configuration snapshots, and many other detailed data. The network-side data also contains data related to the radio coverage prediction that shows the expected coverage in this specific deployment. On the user side, CPEs installed aboard vessels transmitted telemetry that captured device-level QoS metrics (such as RSRP, RSRQ, and cell transitions), GPS-based geolocation, velocity, and diagnostics like speed test results.
The data was highly varied in structure. Some arrived in well-organized formats such as CSV files or relational tables, while other streams came in semi-structured formats like JSON or XML, or as raw, unstructured logs. These were delivered through different channels, including Kafka topics, RESTful APIs, secure file transfers, and compressed archives. The sampling frequency also varied, ranging from high-resolution, near-real-time updates at five-second intervals to lower-frequency aggregated records, depending on the source and purpose of the data. In total, the platform ingested more than 1.5 TB of data each month, representing over 1.5 billion individual records. To properly handle this volume and variety, the system employed a set of scalable data engineering processes that included compression, data selection, enrichment, quality checks, and time alignment. These efforts were essential to ensure the consistency and usability of the data, particularly given the complexity of managing multiple vendor systems in O-RAN.
5.3 Data preparation and feature engineering
Before training the LSTM model, considerable attention was given to refining the dataset to include only meaningful operational conditions. One key challenge was the presence of data collected while vessels were outside the agreed service area, which is beyond the network’s designed coverage. These periods, which occur regularly as part of normal vessel navigation, often produce telemetry that does not reflect an active or stable connection to the offshore network. Including such data could negatively affect the learning process by creating misleading associations between poor connectivity and expected behavior.
((a))
((b))
@@ 5.3 Data preparation and feature engineering
Figure 8: (a) Visualisation for the location of the vessel from the data used in the AI model layered on top of the predicted network coverage. (b) Predicted radio coverage map used for data filtering and feature engineering. Signal strength levels range from strong (green) to weak (darker red shades).
The dataset used in this study combines multiple sources, including network telemetry, onboard CPE data (which includes geolocation), and radio coverage predictions. Merging these sources gave us the full context needed to understand where and when each measurement was taken is relevant to the model. This integration was not only important for filtering out irrelevant records but also played a key role in data engineering. It enabled the creation of new features based on the vessel’s location, observed network performance, and expected signal levels from the coverage map. Without this combined view, building a reliable dataset for training the model would have been far more difficult and time-consuming. An example of this alignment is shown in Fig. LABEL:fig:ue_trace, which visualizes vessel location data overlaid on the predicted network coverage map. As seen in the figure, many data points fall outside the intended coverage boundaries, demonstrating the need for location-based filtering  [^1] The latitude and longitude values are removed for the consideration of the clients’ privacy.. To address this, we first applied a filtering step to exclude all records where location data or time period placed the vessel clearly outside the coverage boundaries.
To illustrate how features were engineered from the integrated data, we used measurements reported by the CPEs such as RSRP, RSRQ, latency, and modem status at five-second intervals, along with GPS coordinates. These were matched against a pre-modeled radio coverage layer that defines expected signal levels across the entire offshore windfarm area. As shown in Fig. LABEL:fig:coloured_coverage, the predicted coverage map was derived from RF planning tools and segmented into levels, where Level 1 (green) indicates the highest expected signal strength, and lower levels (shaded toward red) indicate progressively weaker coverage.
A key feature extracted from this integration was a coverage performance consistency check. For each CPE measurement, the current observed RSRP was compared against the average expected RSRP for that specific geolocation. If the vessel was located in a Level 1 (green) zone, then a certain threshold of signal strength was expected. Any significant deviation from this expected performance, even when the vessel is physically inside a strong coverage area, was flagged and included as a new input feature to the anomaly detection model.
Another example, the distance between the vessel and all known cell sites was calculated. This enabled us to extract not only the distance to the currently connected serving cell but also to determine the closest neighboring cell not currently connected to the CPE. This "nearest-non-serving cell" feature is particularly relevant in assessing whether a better handover decision could have been made.
Meanwhile, the handover activity is tracked. Frequent handovers, especially when occurring in low-mobility or stationary scenarios, were considered potential indicators of instability. The frequency and duration of connections to each serving cell were included as dynamic features in the model’s input sequence.
We also introduced a binary feature that indicated whether the vessel was connected to the list of expected best sites in that location (based on the coverage map) at each timestamp. This “connected-to-best-cell-in-zone” flag, combined with the performance consistency check, helped distinguish between anomalies caused by genuine radio conditions and those arising from device-side or transient network behavior. These spatial, temporal, and performance-based features, along with many others not described here because of their large number and proprietary nature, were included in the LSTM model’s training dataset to help it detect and prevent connectivity problems before they happen2. Table 2 presents some of the features we calculated to feed the LSTM model.
Table 2: Summary of some of the engineered and raw features for anomaly detection model
Feature Name Description Type
distance_to_serving_cell Euclidean or Haversine distance from vessel’s GPS location to the currently connected cell site. Continuous
distance_to_nearest_cell Distance to the geographically closest available (but not connected) cell site. Continuous
rsrp_expected_zone Average RSRP value expected at the vessel’s current geolocation based on RF coverage prediction. Continuous
rsrp_deviation_from_expected Difference between the observed RSRP and expected RSRP at the same location. Continuous
connected_to_best_cell_in_zone Boolean flag indicating if the vessel located in the current location connected to any of the best cells available in that zone. Binary (0/1)
handover_count_window Number of cell handovers during a defined time window. Integer
connected_cell_duration Time duration (in seconds) connected to the current serving cell. Continuous
wan_id Identifier for the wide area network interface used by the CPE. Categorical
carrier Mobile network operator currently serving the connection. Categorical
rsrp Measured Reference Signal Received Power. Continuous
sinr Measured Signal-to-Interference-plus-Noise Ratio. Continuous
rsrq Measured Reference Signal Received Quality. Continuous
latency Round-trip latency as experienced by the CPE. Continuous
time Timestamp of the measurement event. Time
cell_and_network_metrics ... Aggregated cell availability, fault indicators, and thousands of additional performance metrics collected from the RAN and core network layers. Mixed
5.4 LSTM model design and training
@@
The development of the LSTM model adopts the manner of offline development and online deployment. The historical data stored by the platform is used for the model’s training and testing. After the trained model meets the training criteria, it will be deployed in the platform as a container to work online. Historical data includes two parts: the prepared and engineered data and corresponding actions. Actions here are operations recorded by the engineers manually when the QoS is lower than the threshold and result in disconnection.
@@
The next anomaly must be forecasted according to the hidden pattern of the historical network and CPE state records, so a reasonable solution is to formalize a sequence prediction task, where multiple discrete records on sequential timestamps will be cascaded as one training sequence. Then these training sequences are labelled by using manual operations. The training sequence is labeled as ‘0’ if there is an action needed, otherwise, the label is ‘1’.
The features in Table 2, are some of the taken features in this study  [^2] This paper presents a complete overview of the AI-driven anomaly detection system, including data handling, platform architecture, model design, and deployment in a live offshore O-RAN environment. While the feature engineering process and input transformation methods are addressed, their coverage is limited due to commercially sensitive content, proprietary engineering practices and client-specific implementations. As a result, detailed result interpretation and feature-level analysis are discussed only at a high level to maintain transparency without disclosing proprietary information.. Due to the feature engineering, we were able to replace some less important data like latitude, and longitude after using them to calculate more important features such as distance from all cells. The recording interval of fed data is 5 seconds and we used two months of records after being stored and then processed in our cloud-native data platform. The data pre-processing process is illustrated in Fig. 9.
@@
First, for the given raw dataset, there are some records missing. So, the process needs to pad these missing values and remove the records that are obviously out of the normal range. Then as a step of de-noising, the data is downsampled to $1/N$ of the initial one. Then assemble every $L$ single feature vector as one training sequence. Lastly, the training sequence is labeled. It is worth mentioning that the criteria of labeling, are the manual actions taken by the engineer, which may lag over the first occurrence of the anomaly of $5$ to $20$ minutes. So, $10$ minutes is taken as the average delay applied to data labeling. The training sequence will be a matrix with size $1\times 250\times L$, and the labels will be binary values. In this use case, after validation, $N=2$ and $L=6$ are reasonable options.
@@
The architecture of the NN model is depicted in Fig. 10, wherein the LSTM module inherits from the LSTM function in Pytorch and the input size equals $1\times 250\times 6$, hidden size equals 2. Two linear layers follow the LSTM module  [^3] A tutorial of LSTM with detailed architecture and mathematical explanation can be found in [34]., and the output is the indicator of taking action or not. The last layer adopts the activation function ’sigmoid’.
@@
For the model’s training, the training/validation dataset is divided according to the 80%/20% rule. The loss function is the binary cross entropy (BEC) [35]. The training batch size is $2048$ and Adam is taken as the optimizer. In the batch sampling process, the weighted batch sampler is used because this dataset is an imbalanced dataset, which means that labels ’$1$’s are far more than labels ’0’s. The above training dataset is used to train the LSTM model, and the test set is used to evaluate the trained model. In the training stage, the training data is shuffled before batch sampling. Fig. 11 shows the model’s accuracy plots in the training and test. The convergence is reached after $100$ epochs.
@@
5.5 Model deployment and operational behaviour
The LSTM-based anomaly detection model was deployed as a containerized microservice within the Boldyn cloud-native data analytics platform. The model was developed and validated in the application layer, then dockerized and integrated into the DevOps pipeline. It is hosted in the processing layer, where it operates as part of the production environment. The microservice receives a continuous stream of real-time telemetry data from multiple sources, including CPE and O-RAN network functions, via the platform’s data streaming layer. The model processes incoming time-series data to identify patterns indicative of modem performance degradation.
Figure 12: Anomaly detection operational behaviour impact compared to the conventional healing systems
Once an anomaly is detected, the system immediately triggers a predefined control action, such as restarting or reconfiguring the modem, using available network APIs. The entire process, from detection to execution, is automated and designed for near real-time responsiveness. Fig. 12 illustrates the expected operational behavior of the system under two conditions. In the top timeline (with AI, using LSTM), modem degradation is detected early (blue segment), and an automatic recovery action is taken (green transition) before a critical failure occurs. In contrast, the lower timeline (without AI, using conventional methods) where modem degradation progresses unnoticed into a full disconnection (red segment), followed by delayed manual recovery after alarm notifications. This diagram illustrate how the deployed system is architected to intervene earlier and avoid service impact through continuous anomaly monitoring and response automation.
5.6 Result discussion and operational effectiveness
The LSTM-based anomaly detection model was validated using live telemetry from an offshore O-RAN deployment involving dual-modem CPE devices operating under dynamic maritime conditions. Table  3 summarizes weekly average network KPIs before and after the model’s integration into the Boldyn cloud-native platform. Following deployment, the system demonstrated substantial gains in service resilience: single-modem disconnections fell from 170 to 20 per week, while dual-modem dropouts decreased from 25 to just 1. Manual network interventions were virtually eliminated, replaced by automated actions. Average response latency to degradations and disconnections shrank from 30 minutes to approximately 10 seconds. In many of these cases, the AI model intervened before full service was lost, helping prevent modem failures that would have previously required human action or resulted in complete UE disconnection. Prior to adopting the LSTM model, anomaly detection was handled by a rule-based conventional system crafted by radio engineers and SMEs from the experience and the observed cases faced before the disconnection. These rules used fixed limits and set conditions that worked for known problems but couldn’t adapt to new or unexpected issues. In real-world use, the system often failed to detect complex or unexpected failures, since creating rules for every possible case is extremely difficult. In contrast, the LSTM model was trained on past data and can recognize different types of failures by learning patterns over time. It was able to spot small signal issues and short-term changes that weren’t programmed in advance, leading to faster and more accurate responses.
Figure 13: Network Performance Before vs. After AI Deployment (Trip A to B). The chart compares two trips on the same route, one from history before AI applied (dotted lines) using rule-based logic, and one after deploying an LSTM model (solid lines). With AI, the system detected issues early and prevented disconnection, while the legacy setup failed to react in time, leading to a full drop.
To analyse the system’s results more closely, a route-specific comparison provides additional context. Fig. 13 shows telemetry data from the same vessel journey (Trip A to B) recorded at two different times. The first is from historical data, before the LSTM model was deployed, when only the conventional rule-based method was used. The second is from after the LSTM model was deployed. The data covers an 8-minute window and is plotted using dotted lines for the conventional system and solid lines for the LSTM-based system. In the LSTM-enabled case, the model detects the anomaly early (marked by the blue vertical line) and responds quickly, stabilising both RSRP, RSRQ and throughput. In contrast, the conventional system fails to detect the issue, which leads to a full disconnect. If the data from the conventional case had been processed by the LSTM model, it would have triggered an action earlier (marked by the black vertical dotted line). This highlights how the LSTM model performs better at detecting issues compared to the traditional system.
The last two charts in Fig. 13 show that looking only at network-side data or only at CPE data is not enough to understand the full picture. In this case, the network was working well, and there were no issues inside the vessel’s local system. Also, the connectivity chart shows that the device was not always connected to the best cell most of the time. This is expected during vessel movement, especially between coverage sectors and outside the line-of-sight of the cell. In maritime environments, signals can reflect off the water or other surfaces, and sometimes the best signal comes from a tower that is not directly in front of the vessel.
These results (see footnote2) show how complex the O-RAN environment can be. The conditions change often, coverage shifts from place to place, and many factors outside the system can affect performance. The LSTM model works well in this setting because it takes in many types of information, such as connection flags, mobility changes, signal levels, distance, throughput, and service availability and many more. By learning from this combined data, it can notice small changes in behavior that other systems would miss. This helps the system make better decisions, especially in large and mixed deployments where one data source alone cannot explain what is happening.